[figure: 9.7 cm]

Noncollinear magnetism on UFe_{x}Al_{12−x} series: ab-initio calculations

C. Cardoso^{a}, T. Gasche^{a,b}, L. M. Sandratskii^{c} and M. Godinho^{a} aDepartamento de Física, Faculdade de Ciências, Universidade de Lisboa, Portugal

bLaboratório de Física, Academia Militar, Amadora, Portugal

cMax-Planck-Institut fr Mikrostrukturphysik Weinberg 2, D-06120 Halle, Germany

The UFe_{x}Al_{12−x} series has been the object of a systematic experimental study in single crystals with different compositions [1]. Several measurements technics were used in Ref. [1] to suggest a magnetic phase diagram of the system. Three different ranges of composition were found to have different magnetic behavior: x < 4 + δ_{1}, 4 + δ_{1} < x < 5 − δ_{2}, and x > 5 − δ_{2}. The UFe_{4}Al_{8} and UFe_{5}Al_{7} compounds belong to two different regions of the phase diagram. DFT calculations and symmetry analysis for these two compounds [2, 3] contributed to the understanding of the origin of the complex magnetic behavior of the series. In order to complete the systematic study of the series we report calculations of the magnetic properties of UFe_{4.5}Al_{7.5}. The latter compound belongs to the third and the most interesting region of the phase diagram with two magnetic phase transitions. The symmetry analysis determines a magnetic structure with two noncollinear sublattices which is different from that proposed in Ref. [1] This is confirmed by the calculations which yield a magnetic moment in fair agreement with experimental data. We also discuss the phase diagram of the series based on both the experimental and the ab-initio studies.

[1] M. Kuznietz et al., Phys. Rev.B, 60 9494 (1999).

[2] C. Cardoso et al., Phys. Rev. B, 65, (2002) 94413-94417

[3] L.M. Sandratskii and J. Kübler, Phys. Rev. B, 60 (1999) R6961.

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : poster

Corresponding author : C. Cardoso

Address for correspondence : Campo Grande, Edifico C8, piso 4 1749-016 Lisbon, Portugal

Email address : m19908@cc.fc.ul.pt